# NeAR: Coupled Neural Asset–Renderer Stack

### v1

@@ B.1 Additional Comparisons
We provide comprehensive visual comparisons to further substantiate the effectiveness of our method. Figures 11 and 12 illustrate additional results for single-image reconstruction under diverse illumination conditions. For single-image relighting with unknown input lighting, we present extended comparisons in Figures 13 and 14. Notably, our method recovers significantly more accurate shadows and specular highlights compared to existing 2D diffusion-based relighting models [54, 16, 23, 52], which often struggle with physical consistency.
@@
We also conduct detailed comparisons against state-of-the-art 3D generation methods capable of producing PBR materials [60, 5, 3, 6]. As shown in Fig. 15, our approach demonstrates superior material disentanglement, yielding highlights and tonal values that align closely with the ground truth. To ensure a fair comparison and isolate material quality from geometric failures, we provide the baselines with a fixed frontal view and evaluate the rendered output from the same perspective. This setup mitigates the potential for geometric collapse or severe artifacts in baseline methods, focusing the evaluation on rendering and relighting fidelity.
@@ Appendix C Limitations and Future Work
First, the reconstruction of high-frequency details is bottlenecked by the bottleneck of feature compression. The semantic features extracted by DINOv2 [31] undergo significant downsampling and compression via the VAE encoder to fit within the constrained resolution of the LH-SLAT voxel grid. This process inevitably results in information loss, which is particularly evident in objects with intricate textures, such as the blurred text on the magazine cover shown in Fig. 10 (Left). Future work will explore higher-resolution grids or multi-scale feature refinement strategies to better preserve fine-grained details.

### v2

@@ B.1 Additional Comparisons
We provide comprehensive visual comparisons to further substantiate the effectiveness of our method. Figures 14 and 15 illustrate additional results for single-image reconstruction under diverse illumination conditions. For single-image relighting with unknown input lighting, we present extended comparisons in Figures 16 and 17. Notably, our method recovers significantly more accurate shadows and specular highlights compared to existing 2D diffusion-based relighting models [rgbx, jin2024neural, diffusionrenderer, zeng2024dilightnet], which often struggle with physical consistency.
@@
We also conduct detailed comparisons against state-of-the-art 3D generation methods capable of producing PBR materials [zhao2025hunyuan3d, chen20243dtopia, sf3d, chen2025meshgen]. As shown in Fig. 18, our approach demonstrates superior material disentanglement, yielding highlights and tonal values that align closely with the ground truth. To ensure a fair comparison and isolate material quality from geometric failures, we provide the baselines with a fixed frontal view and evaluate the rendered output from the same perspective. This setup mitigates the potential for geometric collapse or severe artifacts in baseline methods, focusing the evaluation on rendering and relighting fidelity.
@@ Appendix C Limitations and Future Work
First, the reconstruction of high-frequency details is hindered by the feature compression step inherent to the pipeline. Specifically, the semantic features extracted by DINOv2 [oquab2023dinov2] undergo substantial downsampling and compression via the VAE encoder to match the limited resolution of the LH-SLAT voxel grid. This compression process inevitably leads to the loss of high-frequency information—a shortcoming that is particularly pronounced for objects with intricate textures, such as the blurred text on the magazine cover illustrated in Fig. 13 (Left). To mitigate this issue, future work will investigate the use of higher-resolution voxel grids or multi-scale feature refinement strategies, both of which aim to better preserve fine-grained details during reconstruction.